# Enhancing Few-shot Keyword Spotting Performance through Pre-Trained Self-supervised Speech Models

### v1

@@ IV Conclusion
This work introduced a novel few-shot keyword spotting framework that leverages pretrained self-supervised transformer models to enable robust keyword detection on resource-constrained edge devices, By training an intermediate teacher model with Sub-center ArcFace loss, and a lightweight student model using knowledge distillation, our approach achieves significant improvements in acoustic separation, even for unseen classes and under cross-domain conditions. Future work will focus on extending this approach to multilingual settings, exploring alternative SSL architectures, and optimizing the deployment on ultra-low-power hardware.
Acknowledgment
We would like to thank Batuhan Gundogdu and Gorkem Ulkar for their valuable contributions to shape this paper.

### v2

@@ IV Conclusion
This work introduced a novel few-shot keyword spotting framework that leverages pretrained self-supervised transformer models for robust detection on resource-constrained edge devices. By training an intermediate teacher model with Sub-center ArcFace loss, and a lightweight student model using knowledge distillation, our approach achieves significant improvements in inter-class discrimination in the embedding space, even for unseen classes and under cross-domain conditions. Future work will focus on extending this approach to multilingual settings, exploring alternative SSL architectures, and optimizing the deployment on ultra-low-power hardware.